OpenStudy (anonymous):

A savings bond offers intrest at rate of 6.6% compunded semi-annually, and a 500$ bond is purchased. Write an equation for the value of the investment as a function of time, in years.

OpenStudy (anonymous):

the general form is A = P (1 + r/n)^nt, where t is years, n is number of compounds per year.

OpenStudy (anonymous):

So plug in r = 0.066, n=2, and P=500; that's your formula.

OpenStudy (anonymous):

but wouldn't I use a=p(1+i)^2t?

OpenStudy (anonymous):

It's the same equation as long as you select i correctly.

OpenStudy (anonymous):

Remember, you're getting the equivalent of 6.6% annually, not 6.6% every compounding period.

OpenStudy (anonymous):

So you divide the annual interest rate by 2, and compound twice as often (which is why the exponent is 2t instead of t)

OpenStudy (anonymous):

It's semi-annually so does that means i do 0.066/2 =i?

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

P (1 + r/n)^nt turns into that for n = 2.

OpenStudy (anonymous):

:) Thank you, it's the first time I'm using this site. Really helpful. Appreciate your help.

OpenStudy (anonymous):

no problem.

OpenStudy (anonymous):

Please click 'good answer' to give medals out to people who help you. :-)

OpenStudy (anonymous):

:) Okie dokie.

OpenStudy (anonymous):

a=p(1+i)^2t a=500(1.033)^2t Is that right?

OpenStudy (anonymous):

yep i think so